OpenStudy (anonymous):

\[\sqrt{6x-17}=4-x\]

OpenStudy (anonymous):

im soooooooo lost

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

it should be clear what you get on the left, because when you square the square root of something, you just get the something

OpenStudy (anonymous):

so if you square both sides you get \(6x-17\) on the left. on the right you get \((4-x)^2=(4-x)(4-x)=16-8x+x^2\)

OpenStudy (anonymous):

now your job is to solve \[6x-17=16-8x+x^2\] which is a quadratic equation. do you know how to solve it?

OpenStudy (anonymous):

no im flipping lost

OpenStudy (anonymous):

if so then you are in good shape. if not, you are stuck put everything on one side of the equal sign to get \[x^2-14x+33=0\]

OpenStudy (anonymous):

hold the phone were the first steps clear or were they confusing?

OpenStudy (anonymous):

no im confused this pellet is killing my brain

OpenStudy (anonymous):

lets start at the beginning. this is what you started with \[\sqrt{6x-17}=4-x\]

OpenStudy (anonymous):

okay so there is no solution yes that is the equation

OpenStudy (anonymous):

you have to eliminate the square roots

OpenStudy (anonymous):

and you do that by squaring both sides

OpenStudy (anonymous):

when you square \[\sqrt{6x-17}\] you get \[6x-17\] right?

OpenStudy (anonymous):

like if i square \(\sqrt{5}\) i get 5

OpenStudy (anonymous):

the only algebra you need to do at this step is square the right hand side, which was \[4-x\] and when you compute \[(4-x)^2\] you get \[(4-x)(4-x)=16-8x+x^2\] is that confusing or is it ok?

OpenStudy (anonymous):

there is a solution you have to solve \[6x-17=16-8x+x^2\] or \[x^2-14x+33=0\] \[(x-3)(x-11)=0\] \[x=3\] or \[x=11\] so those are the possible answers. now you have to check them

OpenStudy (anonymous):

when you do you will see that 11 does not work, but that 3 does

OpenStudy (anonymous):

well im still lost im sorry im mathematically challenged

OpenStudy (anonymous):

go slow don't try to do it all at once this one take several steps, but i think i have written them all out